OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (1, 9); Slope: 5 (Points : 1) y – 1 = 5(x + 9) y – 9 = 5(x + 1) y – 9 = 5(x – 1) y + 9 = 5(x – 1)

OpenStudy (anonymous):

Hey @daisyluv . General equation for point -slope form is y-y1=m*(x-x1) where y1 is the y coordinate of the line and x1is the x coordinate of the line m is the slope of the line putting the values in the equation we get: y-9=5(x-1) -> Option C

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome